OpenStudy (kj4uts):

Which of the following are geometric sequences? A. 1, 3, 9, 27, 81 B. 10, 5, 2.5, 1.25, 0.625, 0.3125 C. 3, 6, 9, 12, 15, 18 D. 5, 10, 20, 40, 80, 160 I think the answers are A, B and D? How do I know?

jimthompson5910 (jim_thompson5910):

let's focus on 1, 3, 9, 27, 81

jimthompson5910 (jim_thompson5910):

divide each term by the previous term 3/1 = 3 9/3 = 3 27/9 = 3 81/27 = 3 Since we get the same result each time, this shows us that the common ratio is 3. This means sequence A is indeed a geometric sequence.